OpenStudy (anonymous):

Tank A is conical in shape with a circular cross section. The diameter of tank A is 0.05 m at its base. The diameter increases with height such that all sides of the tank slope out at a ratio of 1:10 (1 horizontal:10 vertical). (for example at a height of 2 m the tank has a diameter of 0.45 m) Can some one please explain this, also what is the Formula used to solve this?

OpenStudy (anonymous):

@satellite73 would you know?

OpenStudy (anonymous):

@mathslover

OpenStudy (amistre64):

you have not posted a question that asks to be solved ...

mathslover (mathslover):

Amistre has a point :D

OpenStudy (anonymous):

@amistre64 has a point, what is the formula for a conical cylinder? to find the volume.

OpenStudy (anonymous):

volume of a conical cylinder is (1/3)*pi*(r^2)*height

OpenStudy (anonymous):

what class is this from?

OpenStudy (phi):

 the volume of your cone will be the entire cone minus the small cone at the bottom. Using the numbers given, the radius of the tiny cone is 0.05/2 = 0.025 Its height will be 10 times bigger or 0.25

OpenStudy (anonymous):

Problem Solving 4 - I have 3 subjects left till I graduate as an Engineer in Australia. Thanks, I could not get a straight answer from google.

OpenStudy (anonymous):

Thanks phi great explanation, I understand now.

OpenStudy (phi):

You have to be careful of the details. the height of the *entire cone* (including the tiny cone) is h+0.25 the radius at height h will be (h+0.25)/10 as a check, when h=2, r = 2.25/10= 0.225 and the diameter will be 2*0.225= 0.45 which matches with their (for example at a height of 2 m the tank has a diameter of 0.45 m)

OpenStudy (anonymous):

thanks I wish I had another medal.

OpenStudy (phi):

post what you get for the volume of the truncated cone as a function of height h.

OpenStudy (anonymous):

i am making a model in Simulink this is for two tanks which are joined which support the movement of a piston.

OpenStudy (anonymous):

(1/3)*pi*(r^2)*height rearranged for height = Vol / (1/3)*pi*(r^2)

OpenStudy (phi):

You could also use this formula for the volume of a truncated cone http://keisan.casio.com/has10/SpecExec.cgi?id=system/2006/1223372110

OpenStudy (anonymous):

that right phi?

OpenStudy (phi):

yes, though I would write 1/(1/3) as 3, so but we have to be careful, because we have a truncated cone

OpenStudy (anonymous):

ok thanks.

OpenStudy (phi):

if we let H be the height of the entire cone, and h be the height of the truncated cone, you can say H= h+0.25 the volume of the entire cone is \[ V_{\text{entire}} = \frac{\pi}{3} r^2 H \] the volume of the truncated cone is the entire volume minus the little cone \[ V_{\text{truncated}} = \frac{\pi}{3} r^2 H - \frac{\pi}{3} 0.025^2 \cdot 0.25 \] using the fact that r= H/10, we can write \[ V_{\text{truncated}} = \frac{\pi}{3} \frac{H^2}{10^2} H - \frac{\pi}{3} \frac{0.25^2}{10^2} \cdot 0.25 \\ V_{\text{truncated}} = \frac{\pi H^3}{300} - \frac{\pi 0.25^3}{300} \]

OpenStudy (phi):

you can solve for H, and then h= H-0.25